OpenStudy (meehan98):

So, I need some help with Logarithms. The equation is: log (subscript 3) x^2 + log (subscript 3) 9 = 4. I thought you multiply (x^2)(9) in order to get one log on the left side, so you're left with... log (subscript 3) 9x^2=4 and because the log is the exponent, you have log (subscript 3) 9x^2=3^4=81, which makes x=3, but that is the wrong answer. Thank You!

OpenStudy (freckles):

\[\log_3(x^2)+\log_3(9)=4 \\ \text{ yes you can use product rule } \\ \log_3(9x^2)=4\]

OpenStudy (freckles):

you can then write in the equivalent exponential form

OpenStudy (freckles):

\[\log_b(x)=y \implies b^y=x\]

OpenStudy (freckles):

\[3^{4}=9x^2\]

OpenStudy (freckles):

ok and I think you said that above

OpenStudy (freckles):

\[81=9x^2 \\ \frac{81}{9}=x^2 \\ x^2=9\] you should get two solutions not just one

OpenStudy (freckles):

x=3 is right but there is one more solution

OpenStudy (meehan98):

Oh, so it would be 3, -3.

OpenStudy (freckles):

yep

OpenStudy (meehan98):

Ok, thank you! For some reason, I had the mindset that it couldn't be a negative number. You helped so much!

OpenStudy (freckles):

when you plug in -3 into the equation you have that one term is log((-3)^2)) the inside of that log is positive so it is cool

OpenStudy (freckles):

you know since (-3)^2 =9 which is a positive number

OpenStudy (freckles):

your aim is to make sure the solutions you get do not make the inside of the logs negative or zero

OpenStudy (meehan98):

Okay, that makes sense! Thank you!!

OpenStudy (freckles):

\[f(x)=\log(x^2) \text{ has domain all real numbers } \\ \text{ while } g(x)=\log(x) \text{ has domain } x>0\]

OpenStudy (freckles):

so if the equation was: \[2 \log_3(x)+\log_3(9)=4 \\ \text{ you would want \to make sure your solutions satisfy } x>0\]

OpenStudy (freckles):

\[\log_3(x^2)+\log_3(9)=4 \text{ by power rule } \\ \log_3(9x^2)=4 \text{ by product rule } \\ 3^{4}=9x^2 \text{ equivalent exponential form } \\ 81=9x^2 \\ x^2=9 \\ x=\pm 3 \\ \text{ but we said } x>0 \text{ so only } x=3 \\ \text{ works for } 2 \log_3(x)+\log_3(9)=4\]

OpenStudy (freckles):

I made a mistake above with the domain of f

OpenStudy (freckles):

\[f(x)=\log(x^2) \text{ has domain all real numbers excluding } x=0\]

OpenStudy (freckles):

anyways I hope you see I was solving a slightly different equation right there

OpenStudy (freckles):

\[2 \log_3(x)+\log_3(9)=4 \text{ has solutions } x=3 \\ \log_3(x^2)+\log_3(9) =4 \text{ has solutions } x=3,-3\]

OpenStudy (meehan98):

Okay, you have made this a lot easier! The pieces of information finally fit now!! Thank you so much!

OpenStudy (freckles):

np

OpenStudy (freckles):

I gave you medal because you are awesome. You showed your work. And that totally rules in my opinion. You also did pretty well on solving it.

OpenStudy (meehan98):

Thanks! I'm pretty new to this site, and I will definitely stick with it because the help is awesome!